------------------------------------------------------------ Date: Thu Feb 26 07:52:10 2004, Shift: 00:00-08:00 Supervisor: EBJ, with: Log book#15 pages 67 - 67 Shift Summary: ------------------------------------------------------------ 1. The machine had a number of problems today, but we are running now. 2. I continued taking data at 1/2 field A for the FS and full field for the MRS. 3. There were no detector problems. ------------------------------------------------------------ _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Thu Feb 26 07:53:06 2004
This archive was generated by hypermail 2.1.8 : Thu Feb 26 2004 - 07:53:26 EST